Allied Biscuit Co. forecasts the following income statement for the next year:

Income Statement For the Year Ended on December 31

Net sales $960,000
Less: Variable costs 532,000
Less: Fixed costs 151,000
EBIT, or NOI $277,000
Less: Interest 161,000
EBT $116,000
Less: Taxes 46,400
Net income $69,600

Allied Biscuit Co. uses no preferred stock in its capital structure. Click on the following lines and calculate the degrees of operating, financial, and total leverage for Allied Biscuit Co.

DOL (1.59, 1.67,1.75, 1.55,1.42)
DFL (2.12, 2.28, 2.93, 2.56,1.98, 2.74,2.39)
DTL (4.00, 3.63, 3.45, 3.89, 3.79, 3.69,3.54
Answer #1

Degree of Operating leverage (DOL) = ( Sales - Variable cost) / Net Opearating income (NOI)

= ( 9,60,000 - 5,32,000) / 2,77,000

= 4,28,000 / 2,77,000

= 1.545126  

= 1.55 (Rounding of to two decimals)

Degree of Financial leverage (DFL) = Earnings before interest and Taxes (EBIT) / Earnings before Taxes (EBT)

= 2,77,000 / 1,16,000

= 2.387931

= 2.39 (Rounding of to two decimals)

Degree of Total leverage (DTL) = Degree of Operating leverage (DOL) * Degree Financial leverage (DFL)

= 1.545126 * 2.387931 {Taking before rounding up values, for getting accurate Degree of Total leverage (DTL)}

= 3.689655

= 3.68 (Rounding of to two decimals)

Please comment, if any further assitance required.

Thank you.....